QUESTION 10

  1. Power, P, is defined as the rate at which work, W, is done, or the work divided by the time, t, over which that work occurs: P = W / t. Starting from rest, at the bottom of a flight of stairs, what is the power output for a person of mass 55 kg who runs up a flight of stairs of height 3.0 m and reaches the top of the stairs, in a time of 3.0 s, with a final speed of 2.0 m/s?

    a.

    493 J

    b.

    502 J

    c.

    576 J

    d.

    538 J
Answer #1

Solution:

work done = mgh + 0.5 mv^2

=> w = 55 * 9.8 * 3 + 0.5 * 55 * 2 * 2

=> w = 1727 J

Power = w/t

P = 1727 / 3

=> P = 576 W

option C is right.

please comment for queries.

please rate thanks.